OpenStudy (anonymous):

Factor completely. If a polynomial is prime, state this. 9 + 8t2 − 18t

jhonyy9 (jhonyy9):

8t2 -18t +9 =0 so you have wrote this exercise correct sure ?

jhonyy9 (jhonyy9):

t1,2=(18+/- sqrt(324-288))/16 =(18+/-6)/16 = 12/16 =3/4 and 24/16 =3/2 than factorizing this equation we will get (t-3/4)(t-3/2)=(4t-3)(2t-3)